OpenStudy (anonymous):

In coming to rest, suppose that a CD makes one half as many revolutions in a second. How many revolutions does the CD make in coming to rest if it makes 2.6 revolutions in the first second after the stop function is activated?

OpenStudy (theeric):

Hi! Out of curiosity, what subject is this, @michelle.k ?

OpenStudy (anonymous):

algebra 2

OpenStudy (theeric):

Okay, thanks! Have you tried to start this problem yet?

OpenStudy (anonymous):

no. I don't really understand it to be honest

OpenStudy (theeric):

I'm working on it :) The first step is to understand the situation. The second step is to catch it in math.

OpenStudy (anonymous):

OKAY

OpenStudy (theeric):

I don't think my brain is functioning well enough for this problem! Keep bumping it!

OpenStudy (theeric):

Have you done logarithms, yet?

OpenStudy (anonymous):

Seems to me this would run out as follows...2.6 in the first second, then 1.3 in the next second, then .65 in the next second and so on as it would decrease by 1/2 from the previous second. Theoretically it would never stop. Have to work on the formula for this.

OpenStudy (anonymous):

it is a geometric series.

OpenStudy (anonymous):

Since we know \[ a_{n+1}=\frac 12a_{n} \]

OpenStudy (anonymous):

We also know that \(a_1=2.6\).

OpenStudy (anonymous):

Even though technically the CD never comes to a stop, there are a certain number of revolutions it can make even given an infinite amount of time. You need to remember the formula for infinite sums of geometric series.

OpenStudy (anonymous):

Geometric series come in the form \[ a_n = a_1r^{n-1} \]And we know \(a_1=2.6\) and \(r=1/2\).

OpenStudy (anonymous):

OK SO THE ANSWER WOULD BE?

OpenStudy (anonymous):

Are you kidding me? Everyone on this here has become a little more stupid, because you asked for a straight answer after being given an explanation.

OpenStudy (anonymous):

Try asking another question... such as whatever is stopping you from getting the answer yourself.

OpenStudy (anonymous):

IM SO SORRY... I HAD TO USE THE RESTROOM AND WHEN I CAME BACK I HAD ALL THESE EXPLANTATIONS I DIDNT KNOW. I APOLOGIZE

OpenStudy (anonymous):

DO YOU MIND BREAKING IT DOWN FOR ME AGAIN ? I WASNT AT THE COMPUTER SCREEN. my stomach was irritating me so I had to use the bathroom

OpenStudy (anonymous):

Reread what I said. It is still there.

OpenStudy (anonymous):

wouldnt r be 1/2?

OpenStudy (anonymous):

I said that.

OpenStudy (anonymous):

Do you know the equation then for infinite geometric sequence?

OpenStudy (anonymous):

yes s= a/1-r